# On the generalized circular projected Cauchy distribution

### 10 Mar 2026 v2

@@ 2.6.3 Equality of two location parameters
To test the equality of two location parameters, without assuming equality of the parameters $\gamma$ and $\lambda$, we will perform a log-likelihood ratio test [^1] The function to perform the log-likelihood ratio test is available in the R package Directional (Tsagris et al.,, 2026). following (Tsagris et al.,, 2025). Assume we have circular observations from two independent samples, $(\theta_{11},\ldots,\theta_{1{n_{1}}})$ and $(\theta_{11},\ldots,\theta_{1{n_{2}}})$, where $n_{1}$ and $n_{2}$ denote the sample sizes of the two samples.
@@
where $\tilde{a}_{ji}=e^{\tilde{\gamma}_{j}}\cos{\left(\theta_{ji}-\tilde{\omega}\right)}$ and $\tilde{b}_{ji}=\cos^{2}{\left(\theta_{ji}-\tilde{\omega}\right)}+\frac{\sin^{2}\left(\theta_{ji}-\tilde{\omega}\right)}{\tilde{\lambda}_{j}}$, for $j=1,2$.
@@
where $\hat{a}_{ji}=e^{\hat{\gamma}_{j}}\cos{\left(\theta_{ji}-\hat{\omega}_{j}\right)}$ and $\hat{b}_{ji}=\cos^{2}{\left(\theta_{ji}-\hat{\omega}_{j}\right)}+\frac{\sin^{2}\left(\theta_{ji}-\hat{\omega}_{j}\right)}{\hat{\lambda}_{j}}$, for $j=1,2$.
@@ 2.7 Regression modelling revisited
Tsagris and Alzeley, (2025) defined the GCPC regression by considering the bivariate representation (Euclidean coordinates) of the circular data as opposed to their univariate nature (Kato et al.,, 2008). This is akin to the approach employed in the SPML model, as detailed by Presnell et al., (1998) and allows for varying concentration parameter ($\lambda$). The log-likelihood of the GCPC regression model is written as
@@
$\bm{y}_{i}^{\top}\bm{\Sigma}^{-1}_{i}\left(\bm{B}\right)\bm{y}_{i}$ $=$ $\xi_{1i}\left(\bm{B}\right)\xi_{1i}\left(\bm{B}\right)^{\top}/\lambda+\xi_{2i}\left(\bm{B}\right)\xi_{2i}\left(\bm{B}\right)^{\top}$
$=$ $y_{1i}^{2}\left(\xi_{1i}^{2}/\lambda+\xi_{2i}^{2}\right)+y_{2i}^{2}\left(\xi_{2i}^{2}/\lambda+\xi_{1i}^{2}\right)+2y_{1i}y_{2i}\xi_{1i}\xi_{2i}\left(1-1/\lambda\right)$
@@
where $\bm{x}_{i}$ denoting the $i$-th observation of the covariate vector.
To maximize the log-likelihood of the GCPC regression model (20) suggested to use multiple starting points, but we propose a faster option. We begin with initial regression coefficients produced by the SPML regression model of Presnell et al., (1998) and then estimate the $\lambda$ parameter using Brent’s algorithm [^2] This is available in R via the buit-in optimize() function. (Brent,, 2013). Then use these as starting points in R’s built-in optim() function.
Tsagris and Alzeley, (2025) did not consider the case of circular-circular regreression, where a covariate $X$ is circular as well. But, this case is easily adopted in the above case scenario by transforming the circular covariate to Euclidean by using $\bm{z}_{i}=\left(\cos{x_{i}},\sin{x_{i}}\right)^{\top}$. In case of multiple circular covariates, all of them are transformed into their Euclidean coordinates and the same link function (21) is used. Simplicial predictors (compositional data) are straighforward to add by using two options. The first is to apply the additive log-ratio transformation (Aitchison,, 2003), so that the regression coefficients sum to zero and have a nice (derivative based) interpretation. In case of zero values present the logarithmic transformation breaks down, so the alternative approach is to transform them via the $\alpha$–transformation (Tsagris et al.,, 2011). The second approach bypasses the problem of zero values but lacks interpretation of the estimated regression coefficients and requires to choose the value of the $\alpha$ parameter. The choice of $\alpha$ may be overcome by choosing a small value of $\alpha$ (say 0.01) to resemble the isometric log-ratio transformation (Egozcue et al.,, 2003).
An extra extension would be to link the $\lambda$ parameter to the covariates as well, but that would increase the computational cost. Another evident approach to make the model non-linear would be to use splines in the covariates as well.
@@ 3 Simulation studies
We performed simulation studies to examine the performance of the two sample means test and also the effect of the $\lambda$ parameter on the performance of the log-likelihood test.
@@ 3.1 One location parameter
We selected six sample sizes, and generated values from the $GCPC\left(2,3,2\right)$. For each sample size we computed the type I error (assuming $\alpha=0.05$) based on 1,000 replicates. Table 1 presents the results, where we can see that the test slightly overestimates the nominal type I error, but not significantly.
@@ 3.2 Two location parameters
We compared the test when both samples come from the GCPC distribution and when assuming that they come from the CIPC distribution ($\lambda=1$). Six pairs of unequal sample sizes were used, where the true angular mean was $\omega=2$, but the $\gamma$ and $\lambda$ parameters differed. The parameters for the smaller sample were equal to $\gamma=4$ and $\lambda=1$, while for the larger sample they were equal to $\gamma=2$ and $\lambda=3$. Circular data were generated from two GCPC distributions with the specified parameters and the two log-likelihood ratio tests were performed. This process was repeated 1,000 times and the proportion of rejection of the $H_{0}$ (at the $\alpha=0.05$ level) served as an estimate of the type I error. Table 2 presents the results. The GCPC based log-likelihood ratio always attained the correct size, whereas the CIPC based log-likelihood ratio test overestimated the size of the test.
@@ 3.3 Empirical asymptotics for the regression model
where $\bm{u}$ was generated from a von Mises distribution with mean equal to 2 and concentration parameter equal to 5. For the continuous predictor, the model was
@@
Table 3: Matrix $\bm{B}$ of the regression coefficients used in the simulations. The first two columns correspond to the coefficients of the model with the circular predictor, the second two to the model with the continuous predictor and the last two to the model with both of them. In the last case $\bm{\beta}_{1}$ and $\bm{\beta}_{2}$ refer to the circular predictor and $\bm{\beta}_{3}$ to the continous predictor.
@@
Figure 2: Average squared Frobenius norm $\bar{F}^{2}$ versus sample size, both in log-scale for the circular, the continuous and the combination of both.
@@ 4 Real data analysis
We used the same dataset used in Tsagris and Alzeley, (2025), the speed.wind2 data that is available in the R package NPCirc (Oliveira et al.,, 2014). This dataset consists of 199 hourly observations of wind direction and wind speed in winter season (from November to February) from 2003 until 2012 in the Atlantic coast of Galicia (Spain).
4.1 Maximum likelihood estimation
Table 4 presents the estimated parameters of the CIPC and GCPC models. Application of the log-likelihood ratio test to discriminate between the two models (Tsagris and Alzeley,, 2025) clearly favors the GCPC model. Figure 3(a) visualizes the data, it is the circular plot with the mean directions showed with arrows. Evidently the data look bimodal and the CIPC has chosen one of the two modes, whereas the GCPC has chosen a direction that ”cuts” the data into two. The density plot on Figure 3(b) shows that the GCPC density has a similar pattern to the one produced by the kernel density estimate, in contrast to the CIPC distribution that has failed to fit the data. This is more evident in the circular density plot (Figure 3(c)). The CIPC distribution has a circular shape, whereas the GCPC distribution has an elliptical shape, capturing the shape of the data more accurately. Finally, Figure 3(d) visualizes the 95% confidence interval for the mean direction.
Table 4: Estimated parameters of the CIPC and GCPC models applied to the wind direction data.
@@ 4.1 Maximum likelihood estimation
(c) Circular density plot (d) 95% confidence interval for the mean direction
Figure 3: (a) Circular plot of the speed direction data. (b) Density plot of the kernel density estimate, and the fitted CIPC and GCPC distributions. (c) Circular density plot of the two fitted distributions. The green line is the estimated mean direction based on the CIPC, whereas the blue lines is the estimated mean direction based on the GCPC model.
@@ 4.2 Regression
We spotted a mistake in the regression modelling of Tsagris and Alzeley, (2025), where the continuous predictor is the speed of the wind. Table 5 shows the regression coefficients computed in Tsagris and Alzeley, (2025) and the correct estimates. Note that this time the sign of the slope coefficients matches that of the CIPC and of the spherically projected multivariate linear regression models models. The $\lambda$ coefficient still remains statistically significantly far from 1, and the log-likelihood ratio test still favors the GCPC regression over the CIPC regression model.
@@
Model GCPC (Tsagris and Alzeley,, 2025) GCPC
@@ 5 Conclusions
We derived the relationship between the GCPC and the CIPC (reparameterized wrapped Cauchy) distribution, provided a not so convenient formula for the mean resultant length, an alternative analytical formula to compute probabilities, and an analytical formula for the entropy of the GCPC. We further proposed two log-likelihood ratio tests for hypothesis testing with one or two location parameters. For the two location parameters testing case, the test is size correct when the data are wrongfully assumed to follow the CIPC distribution. Finally we corrected a mistake in the regression setting and proposed a faster altenative than the one proposed by Tsagris and Alzeley, (2025).

### 25 Mar 2026 v3

@@ 2.8.3 Equality of two location parameters
To test the equality of two location parameters, without assuming equality of the parameters $\gamma$ and $\lambda$, we will perform a log-likelihood ratio test [^2] All the relevant functions are available in the R package Directional (Tsagris et al., 2026). following (Tsagris et al., 2025). Assume we have circular observations from two independent samples, $(\theta_{11},\ldots,\theta_{1{n_{1}}})$ and $(\theta_{21},\ldots,\theta_{2{n_{2}}})$, where $n_{1}$ and $n_{2}$ denote the sample sizes of the two samples.
@@
where $\tilde{a}_{ji}=\tilde{\gamma}_{j}\cos{\left(\theta_{ji}-\tilde{\omega}\right)}$ and $\tilde{b}_{ji}=\cos^{2}{\left(\theta_{ji}-\tilde{\omega}\right)}+\frac{\sin^{2}\left(\theta_{ji}-\tilde{\omega}\right)}{\tilde{\lambda}_{j}}$, for $j=1,2$.
@@
where $\hat{a}_{ji}=\hat{\gamma}_{j}\cos{\left(\theta_{ji}-\hat{\omega}_{j}\right)}$ and $\hat{b}_{ji}=\cos^{2}{\left(\theta_{ji}-\hat{\omega}_{j}\right)}+\frac{\sin^{2}\left(\theta_{ji}-\hat{\omega}_{j}\right)}{\hat{\lambda}_{j}}$, for $j=1,2$.
@@ 2.9 Regression modelling revisited
Tsagris and Alzeley (2025) defined the GCPC regression by considering the bivariate representation (Euclidean coordinates) of the circular data as opposed to their univariate nature (Kato et al., 2008). This is akin to the approach employed in the SPML model, as detailed by Presnell et al. (1998) and allows for varying concentration parameter ($\lambda$). The log-likelihood of the GCPC regression model is written as
@@
$\bm{y}_{i}^{\top}\bm{\Sigma}^{-1}_{i}\left(\bm{B}\right)\bm{y}_{i}$ $=$ $\bm{y}_{i}^{\top}\left[\xi_{1i}\left(\bm{B}\right)\xi_{1i}\left(\bm{B}\right)^{\top}/\lambda+\xi_{2i}\left(\bm{B}\right)\xi_{2i}\left(\bm{B}\right)^{\top}\right]\bm{y}_{i}$
$=$ $y_{1i}^{2}\left(\xi_{2i}^{2}/\lambda+\xi_{1i}^{2}\right)+y_{2i}^{2}\left(\xi_{1i}^{2}/\lambda+\xi_{2i}^{2}\right)+2y_{1i}y_{2i}\xi_{1i}\xi_{2i}\left(1-1/\lambda\right)$
$=$ $\left(y_{1i}\xi_{2i}-y_{2i}\xi_{1i}\right)^{2}/\lambda+\left(y_{1i}\xi_{1i}+y_{2i}\xi_{2i}\right)^{2}.$
@@
with $\bm{x}_{i}$ denoting the $i$-th observation of the covariate vector.
To maximize the log-likelihood of the GCPC regression model (21), Tsagris and Alzeley (2025) suggested the use of multiple starting values, however we propose a computationally more efficient approach. We begin with initial regression coefficients produced by the SPML regression model of Presnell et al. (1998) and then estimate the $\lambda$ parameter using Brent’s algorithm [^3] This is available in R via the built-in optimize() function. (Brent, 2013). These estimates are subsequently used as starting values in R’s built-in optim() function.
Tsagris and Alzeley (2025) did not consider the case of circular-circular regression, where a covariate $X$ is circular. But, this case is easily accommodated in the above case scenario by transforming the circular covariate to Euclidean by using $\bm{z}_{i}=\left(\cos{x_{i}},\sin{x_{i}}\right)^{\top}$. In case of multiple circular covariates, all of them are transformed into their Euclidean coordinates and the same link function (22) is used. Simplicial predictors (compositional data) are straightforward to add by using two options. The first is to apply the additive log-ratio transformation (Aitchison, 2003), so that the regression coefficients sum to zero and have a meaningful (derivative based) interpretation. In case of zero values present the logarithmic transformation breaks down, so the alternative approach is to transform them via the $\alpha$–transformation (Tsagris et al., 2011). The second approach bypasses the problem of zero values but lacks interpretation of the estimated regression coefficients and requires one to select the value of the $\alpha$ parameter. The choice of $\alpha$ may be overcome by choosing a small value of $\alpha$ (for instance, $\alpha=0.01$) to resemble the isometric log-ratio transformation (Egozcue et al., 2003).
A further extension consists of linking the $\lambda$ parameter to the covariates, but that would increase the computational cost. A complementary approach to introduce non-linearity consists of incorporating splines for the covariate effects.
@@ 3 Simulation studies
We performed simulation studies to examine the performance of the test for one and two location parameters and to assess the effect of misspecifying $\lambda$ on the size of the log-likelihood ratio test. We did not perform simulation studies for the asymptotic confidence interval for the location parameter since this is derived from the log-likelihood ratio test.
@@ 3.1 One location parameter
We selected six sample sizes, and generated values from the $GCPC\left(2,3,2\right)$. For each sample size we computed the type I error (assuming $\alpha=0.05$) based on 1,000 replicates. Table 1 presents the results, where we can observe that the test slightly overestimates the nominal type I error, though the deviation is not substantial.
@@ 3.2 Two location parameters
We compared the test when both samples come from the GCPC distribution and when assuming that they come from the CIPC distribution ($\lambda=1$). Six pairs of unequal sample sizes were used, where the true location parameters were equal to $\omega=2$ for both populations, but the $\gamma$ and $\lambda$ parameters differed between the two populations. The parameters for the smaller sample were equal to $\gamma=4$ and $\lambda=1$, while for the larger sample they were equal to $\gamma=2$ and $\lambda=3$. Circular data were generated from two GCPC distributions with the specified parameters and the two log-likelihood ratio tests were performed. This process was repeated 1,000 times and the proportion of rejection of the $H_{0}$ (at the $\alpha=0.05$ level) served as an estimate of the type I error. Table 2 presents the results. The GCPC based log-likelihood ratio always attained the correct size, whereas the CIPC based log-likelihood ratio test overestimated the size of the test.
@@ 3.3 Empirical asymptotics for the regression model
where $\bm{u}$ was generated from a von Mises distribution with location parameter $\omega=2$ and concentration parameter $\gamma=5$. For the continuous predictor, the model was
@@
Table 3: Matrix $\bm{B}$ of the regression coefficients used in the simulations. The first two columns correspond to the coefficients of the model with the circular predictor, the second two to the model with the continuous predictor and the last two to the model with both of them. In the last case $\bm{\beta}_{1}$ and $\bm{\beta}_{2}$ refer to the circular predictor and $\bm{\beta}_{3}$ to the continuous predictor.
@@
Figure 2: Average squared Frobenius norm $\bar{F}^{2}$ versus sample size, both in log-scale for the circular predictor, continuous predictor and the combination of both.
@@ 4 Real data analysis
We used the same dataset used in Tsagris and Alzeley (2025), the speed.wind2 data that is available in the R package NPCirc (Oliveira et al., 2014). This dataset consists of 199 hourly observations of wind direction and wind speed in winter season (from November to February) from 2003 until 2012 in the Atlantic coast of Galicia (Spain).
4.1 MLE
Table 4 presents the MLE of the CIPC and the GCPC distributions fitted to the speed direction data. We highlight that the MLE using Eq. (7), Tsagris and Alzeley (2025) was trapped in one of the three local maxima, while the MLE using (8) obtained the global maximum. Notice that the estimated $\lambda$ falls within Case D, $\lambda\leq\frac{1}{2}$, and hence the distribution is bimodal, which is also evident from Figures 3(b) and 3(c). Table 4 presents the estimated parameters of the CIPC and GCPC models. Application of the log-likelihood ratio test to discriminate between the two models (Tsagris and Alzeley, 2025) clearly favors the GCPC model.
Figure 3(a) visualizes the data, it is the circular plot with the location parameter shown with arrows. The density plot on Figure 3(b) shows that the GCPC density has a similar pattern to the one produced by the kernel density estimate, in contrast to the CIPC distribution that does not adequately capture the distributional features of the data. This is more evident in the circular density plot (Figure 3(c)). The CIPC distribution has a circular shape, whereas the GCPC distribution has an elliptical shape, capturing the shape of the data more accurately. Figure 3(d) visualizes the 95% confidence interval for the location parameter. Figure 4 visualizes the profile log-likelihood of the location parameter $\omega$. The log-likelihood contains four modes, three of them correspond to local maxima and one is the global maximum.
Table 4: Estimated parameters of the CIPC and GCPC models applied to the wind direction data. The GCPC* indicates the MLE using the polar coordinates parametrized GCPC density (8).
@@ 4.1 MLE
(c) Circular density plot (d) 95% confidence interval for the true location parameter
Figure 3: (a) Circular plot of the speed direction data. (b) Density plot of the kernel density estimate, and the fitted CIPC and GCPC distributions. (c) Circular density plot of the two fitted distributions. The green line is the estimated location parameter based on the CIPC, whereas the blue line is the estimated location parameter based on the GCPC model.
Figure 4: Profile log-likelihood of the location parameter $\omega$ for the GCPC distribution using the speed direction data. The white vertical line corresponds to the estimated location parameter that maximizes the log-likelihood.The periodicity is equal to $\pi/2$.
@@ 4.2 Regression
We identified a mistake in the regression modelling of Tsagris and Alzeley (2025), where the continuous predictor is the speed of the wind. Table 5 shows the regression coefficients computed in Tsagris and Alzeley (2025) and the correct estimates. Note that this time the sign of the slope coefficients matches that of the CIPC and of the spherically projected multivariate linear regression models. The $\lambda$ coefficient still remains statistically significantly far from 1, and the log-likelihood ratio test still favors the GCPC regression over the CIPC regression model.
@@
Model GCPC (Tsagris and Alzeley, 2025) GCPC
@@ 5 Conclusions
We derived the relationship between the GCPC and the CIPC (reparameterized WC) distribution, provided non-analytical formulas for the mean resultant length and the Kullback-Leibler divergence, an alternative analytical formula to compute probabilities, and an analytical formula for the entropy of the GCPC. We also showed the conditions under which the GCPC distribution is unimodal. We further proposed two log-likelihood ratio tests for hypothesis testing with one or two location parameters. For the two location parameters testing case, the test is size correct when the data are wrongfully assumed to follow the CIPC distribution. We observed that in the case of a bimodal distribution there is danger the MLE to be trapped in a local maximum and showed how to escape this trap. Finally, we corrected a mistake in the regression setting and proposed a computationally more efficient alternative than the one proposed by Tsagris and Alzeley (2025).